Round to four decimal places as needed)
Answer #1

Mean = \mu = 177.4

Standard deviation = \sigma = 36.3

Sample size = n = 60

We have to find P( \bar{x} > 145)

For finding this probability we have to find z score.

z =\frac{\bar{x}-\mu }{\sigma /\sqrt{n}}

z =\frac{145-177.4 }{36.3 /\sqrt{60}}=\frac{-32.4}{4.68631}=-6.91

That is we have to find P(Z > - 6.91)

P(Z > - 6.91) = 1 - P(Z < - 6.91) = 1 - 0.000 = 1.0000